# Knowledge Editing for Large Language Models: A Survey

### v3

@@ 5.1. Categorization of KME Methods
The above categorization is achieved based on where (e.g., external parameters or internal weights) and how (e.g., via optimization or direct incorporation) new information is introduced into the LLM during editing. Specifically, methods from each category exhibit different strengths and weaknesses regarding the four crucial evaluation metrics introduced in Sec. 4. For example, external memorization prevails in scenarios that require massive editing while the computational resources are limited, as the size of the memory is controllable to fit into different requirements. On the other hand, global optimization is advantageous when practitioners focus more on the generality of edited knowledge, as the optimization can promote the learning of relevant knowledge (Aghajanyan et al., 2021). The taxonomy is visually illustrated in Fig. 3, and a more detailed demonstration of each category is presented in Fig. 4. The specific characteristics of all methods are summarized in Table 2.
@@ 5.2.2. Memory-based Strategies
In memory-based strategies, the external memory, outside the intrinsic architecture of the pre-trained LLM, functions as a repository to store edited knowledge. The LLM can access and update this memory as required during inference.
@@
In KAFT (Li et al., 2023c), controllability is achieved through the utilization of counterfactual data augmentations. In this approach, the entity representing the answer within the context is substituted with an alternative but still plausible entity. This substitution is intentionally designed to introduce a conflict with the genuine ground truth, thereby enhancing the controllability and robustness of LLMs with respect to their working memory. The aim is to ensure that LLMs remain responsive to pertinent contextual information while filtering out noisy or irrelevant data.
@@ 5.2.3. Extension-based Strategies
Alternatively, a different technique involves integrating an adapter into a specific layer of a pre-trained model. This adapter consists of a discrete dictionary comprising keys and values, where each key represents a cached activation generated by the preceding layer and each corresponding value decodes into the desired model output. This dictionary is systematically updated over time. In line with this concept, GRACE (Hartvigsen et al., 2022) introduces an adapter that enables judicious decisions regarding the utilization of the dictionary for a given input, accomplished via the implementation of a deferral mechanism. It is crucial to achieve a balance between the advantages of preserving the original model’s integrity and the practical considerations associated with storage space when implementing this approach.
@@ 5.3.1. Overview
Different from external memorization-based methods that introduce new parameters to assist the editing of pre-trained LLMs, there also exist branches of works that do not rely on external parameters or memory. Concretely, global optimization strategies aim to inject new knowledge into LLMs by updating all parameters, i.e., $\phi$ in Eq. (13). Through fine-tuning model parameters with specific designs to ensure the preservation of knowledge irrelevant to the target knowledge $t^{*}$, the LLMs are endowed with the ability to absorb new information without altering unedited knowledge. Generally, the goal of global optimization methods can be formulated as follows:
@@
where $f_{\phi}$ denotes the LLM before editing with the pre-trained parameter $\phi$, and $f_{\phi^{*}}$ denotes the edited LLM with updated parameter $\phi^{*}$. Generally, these methods focus more on the precision and generality of desirable knowledge, as the fine-tuning process ensures that the LLMs achieve satisfactory results regarding the edits and relevant knowledge. Nevertheless, as fine-tuning affects all parameters, they cannot easily preserve the locality of edited models, i.e., maintaining consistent output for unedited knowledge (Yao et al., 2023). In practice, directly applying fine-tuning strategies typically exhibits suboptimal performance on KME due to overfitting concerns (Wang et al., 2023b; Meng et al., 2023). Furthermore, fine-tuning large language models is also time-consuming and lacks scalability for multiple edits. Therefore, recently, motivated by these two challenges in fine-tuning, several global optimization works have been proposed and can be categorized as constrained fine-tuning methods and intermediate fine-tuning methods.
@@ 5.3.2. Constrained Fine-tuning
The updates are regularized by restricting the norm of parameters before and after updating. Although restricting the norm is helpful in preventing the forgetting of original knowledge, the fine-tuning process can be less effective. To deal with this, RecAdam (Chen et al., 2020), in addition to the norm constraint, applies an annealing technique to control the ratio between the parameter norm and the fine-tuning loss as follows:
@@
Here $k$ and $t_{0}$ are hyper-parameters. $t$ is the number of fine-tuning steps. Such a design enables a gradual fine-tuning process that prevents massive parameter updates at the beginning. Motivated by the intuition of regularization to preserve original knowledge, PPA (Lee et al., 2022) employs LoRA (Hu et al., 2021) in the feed-forward (FFN) layers of the transformer decoder. LoRA is proposed to train the expansion/reduction matrix, instead of the model parameter $\phi$, to improve training speed by only updating parameters with a low intrinsic rank via dimensionality reduction. PPA leverages plug-in modules trained with constraints via LoRA to keep original knowledge intact. Moreover, the authors assess whether the content of the inputs falls within the scope of $\mathcal{X}_{\mathcal{E}}$ using the K-adapter module (Wang et al., 2020), and redirect such inputs to the new plug-in modules. This information is then used to determine whether to employ LoRA within the FFN layers. To preserve original knowledge, F-Learning (Forgetting before Learning) (Ni et al., 2023) proposes to learn knowledge parameters $\Delta\phi$ that indicates old knowledge to be forgotten, defined as follows:
@@ 5.3.3. Intermediate Fine-tuning Strategies
While constrained fine-tuning techniques have demonstrated remarkable efficacy in a variety of NLP tasks (Wortsman et al., 2022; Ziegler et al., 2019; Bakker et al., 2022), they still exhibit instability and high computational cost when applied to KME, primarily due to the necessity of altering all parameters (Yao et al., 2023). A potential solution to address this challenge is to utilize an intermediate model to obtain the updated parameters in an efficient manner. Such an intermediate model is required to maintain significantly fewer parameters to ensure efficiency (Cheng et al., 2023b). In general, recent works have widely adopted the Hyper-Network (Ha et al., 2016) as the intermediate model. Specifically, the Hyper-Network is a small network that generates the weights for a larger network, referred to as the main network. Specifically, the Hyper-Network takes inputs that contain information about the structure of the weights and generates the weights for layers in the main network. With the generated weights, the main network is updated to map input data to desired output targets. The updating process for the main network, denoted as $\phi$, can be defined as follows:
(23) $\phi^{*}$ $=\phi+\Delta\phi,$
$\Delta\phi$ $=\text{H}(\nabla_{\phi}\mathcal{L}(f_{\phi}(x),y^{*})),\ \text{where}\ x,y^{*}\in\mathcal{X}_{\mathcal{E}},\mathcal{Y}^{*}_{\mathcal{E}},$
@@
However, hyper-networks are generally not capable of updating large language models due to the massive parameter size. To tackle this challenge, MEND (Mitchell et al., 2022a) adopts a mechanism referred to as gradient decomposition. In particular, it leverages small auxiliary editing networks to transform the gradients obtained by standard fine-tuning into edits of weights in a pre-trained model. As gradients are generally high-dimensional objects, a low-rank decomposition of the gradients is utilized to achieve the transformation. Particularly, MEND parameterizes the gradient mapping functions as MLPs with a single hidden layer, such that a significantly small number of parameters are required, compared with the edited models. In this manner, MEND enables fast model editing that can operate on considerably large pre-trained language models. Moreover, KGEditor (Cheng et al., 2023b) proposes to combine the benefits of memory-based methods and hyper-networks to ensure flexibility and further reduce computation costs. Particularly, KGEditor introduces an additional layer with the same architecture of FFN layers for storing knowledge. Then it constructs a hyper-network based on a bi-directional LSTM (Hochreiter and Schmidhuber, 1997) that encodes embeddings of triples. In this manner, KGEditor becomes an efficient way to edit knowledge graph embeddings.
@@ 5.3.4. Summary
KME methods based on global optimization typically apply specific fine-tuning restrictions to regularize parameter updates, namely constrained fine-tuning strategies. This is to prevent overfitting and ensure the model’s performance on the unedited knowledge. One crucial advantage of such strategies is its generality regarding the relevant knowledge, i.e., in-scope inputs $\mathcal{X}_{e}$ of edit $e$. As the global optimization affects all parameters in a language model, the relevant knowledge in it will also be edited, thereby generalizing to such knowledge. On the other hand, the high computation costs of fine-tuning all parameters also motivate researchers to propose intermediate fine-tuning strategies that leverage hyper-networks. Furthermore, global optimization methods are mostly model-agnostic, which means they can be applied to other editing methods. Nevertheless, such possibilities are less explored in the context of KME. In terms of the drawbacks, global optimization methods are suboptimal in maintaining the locality of edited models, as the optimization can easily influence unedited knowledge. Hence, it is crucial to achieve a balance between generality and locality when optimizing language models with specific constraints or intermediate designs.
@@ 5.4.1. Overview
In the following subsections, we first introduce the concept of knowledge neuron in LLMs, which are specific neurons that store factual knowledge and can be activated to generate the desirable answer based on a certain query $x$. Then we discuss two local modification strategies for KME: (1) the gradient-based strategies, which identify and edit knowledge neurons based on gradients calculated from the model outputs; (2) the prompt-based strategies, which locate knowledge neurons based on the input prompts.
@@
(25) $\operatorname{SelfAtt}_{i}(\textbf{x})$ $=\operatorname{Softmax}\left(\textbf{q}_{i}\textbf{k}_{i}^{\top}\right)\textbf{v}_{i},$
$\operatorname{FFN}(\textbf{h})$ $=\operatorname{GELU}\left(\textbf{h}\textbf{W}_{1}\right)\textbf{W}_{2}.$
@@
5.4.2. Gradient-based Strategies
Based on the knowledge neuron view of the FFN layer weights in pre-trained LLMs, various gradient-based methods are proposed to locate and edit the pre-trained LLMs. Generally, they perform editing in a top-down manner, utilizing the supervision signal provided by the correct answer $y^{*}$. As an exemplar work, KD (Dai et al., 2022) proposes to change each weight $w^{(l)}_{i}$ (i.e., the $i$-th weight in the $l$-th layer of FFN) from 0 to the pre-trained value $\hat{w}^{(l)}_{i}$. Then it calculates the cumulative change in the probability of predicting the output $y^{*}$ with input $x$, where the weights with a high cumulative probability are considered relevant for knowledge regarding $y^{*}$. In contrast to locating and editing an individual weight ${w}^{(l)}_{i}$ in the FFN layer, ROME (Meng et al., 2022) proposes to update a whole FFN layer to encode the new knowledge $y^{*}$. Specifically, they view the second layer weights $\textbf{W}_{2}$ in the FFN layer in Eq. (25) as a linear associative memory (Kohonen, 1972; Anderson, 1972) in the form of $\textbf{K}\textbf{W}_{2}=\textbf{V}$, where the keys K and values V associated with $\textbf{W}_{2}$ can be directly calculated with Moore-Penrose pseudoinverse. With such a view of $\textbf{W}_{2}$ in the FFN layer, the optimization objective of updating it into $\hat{\textbf{W}}_{2}$ to encode new knowledge in the edit $e=(s,r,o\rightarrow o^{*})$ can be formulated as follows:
@@ 5.4.2. Gradient-based Strategies
Here $\textbf{k}^{*}$, which should encode the information of the subject $s$, is calculated by sampling multiple $x\sim\mathcal{X}_{e}$ and taking the average of the output of the first dense layer of the FFN. The target activation $\textbf{h}^{*}$ is calculated via optimizing the probability of outputting the correct answers $y^{*}\in\mathcal{Y}_{e}$ of the pre-trained LLM via the subsequent layers. Then, an efficient rank-one update is conducted on the weights $\textbf{W}_{2}$ according to Eq. (26), such that after the update, the edited FFN layer can output the correct hidden representation $\textbf{h}^{*}$ conducive to the generation of the right answer $y^{*}$ from $\textbf{k}^{*}$. Recently, MEMIT (Meng et al., 2023) proposes to further generalize the above editing strategy of the FFN layers of pre-trained LLMs to the mass editing of different knowledge. Particularly, with $u$ new edits $\{e_{1},e_{2},\dotsc,e_{u}\}$ that are required to be updated in the weights $\textbf{W}_{2}$, the mass editing problem can be formulated as the following optimization problem:
@@
where $\textbf{k}_{i}$, $\textbf{v}_{i}$ are the original key, value pairs associated with the weights $\textbf{W}_{2}$ (i.e., row vectors in matrices K, V in Eq. (26)), whereas $\textbf{k}_{i}^{*}$, $\textbf{v}^{*}_{i}$ are the updated key, value pairs calculated from the $i$-th edit $e_{i}$ as with Eq. (26). In addition, since multiple edits are required, the update is shared among different MLP layers, which is conducted in a top-down manner to prevent the potential issue of editing layers that could affect the ones that have already been edited. The residual for each edit is spread evenly over the range of the critical FFN layer. The strategy of residual attribution has recently been improved by PMET (Li et al., 2023b), which adopts a square root strategy to spread residuals to bottom FFN layers such that more precise information can be conveyed to critical layers.

### v4

@@ 5.1. Categorization of KME Methods
The above categorization is achieved based on where (e.g., external parameters or internal weights) and how (e.g., via optimization or direct incorporation) new knowledge is introduced into the LLM during editing. Methods in each category exhibit different strengths and weaknesses regarding the four crucial evaluation metrics introduced in Sec. 4. For example, external memorization prevails in scenarios that require massive editing while the computational resources are limited, as the size of the memory is controllable to fit into different requirements. On the other hand, global optimization is advantageous when practitioners focus more on the generality of edited knowledge, as the optimization can promote the learning of relevant knowledge (Aghajanyan et al., 2021). The taxonomy is visually illustrated in Fig. 3, and a more detailed demonstration of each category is presented in Fig. 4.
@@ 5.2.2. Memory-based Strategies
In memory-based strategies, the external memory, outside the intrinsic architecture of the pre-trained LLM, functions as a repository to store edited knowledge. Here the edits are generally converted to text via pre-defined templates (Zhong et al., 2023; Zheng et al., 2023; Wang et al., 2023a). The LLM can access and update this memory as required during inference.
@@
In KAFT (Li et al., 2023b), controllability is achieved through the utilization of counterfactual data augmentations. In this approach, the entity representing the answer within the context is substituted with an alternative but still plausible entity. This substitution is intentionally designed to introduce a conflict with the genuine ground truth, thereby enhancing the controllability and robustness of LLMs with respect to their working memory. The aim is to ensure that LLMs remain responsive to pertinent contextual information while filtering out noisy or irrelevant data.
@@ 5.2.3. Extension-based Strategies
Alternatively, a different technique involves integrating an adapter into a specific layer of a pre-trained model. This adapter consists of a discrete dictionary comprising keys and values, where each key represents a cached activation generated by the preceding layer and each corresponding value decodes into the desired model output. This dictionary is systematically updated over time. In line with this concept, GRACE (Hartvigsen et al., 2023) introduces an adapter that enables judicious decisions regarding the utilization of the dictionary for a given input, accomplished via the implementation of a deferral mechanism. It is crucial to achieve a balance between the advantages of preserving the original model’s integrity and the practical considerations associated with storage space when implementing this approach. COMEBA-HK (Li et al., 2024a) incorporates hook layers within the neural network architecture. These layers allow for the sequential editing of the model by enabling updates to be applied in batches. This approach facilitates the integration of new knowledge without requiring extensive retraining of the entire model, making it a scalable solution for continuous learning and adaptation. SWEA (Li et al., 2024b) focuses on altering the embeddings of specific subject words within the model. By directly updating these embeddings, the method can inject new factual knowledge into the LLMs. This approach ensures that the updates are precise and relevant, thereby enhancing the model’s ability to reflect current information accurately.
@@ 5.3.1. Overview
Different from external memorization methods that introduce new parameters to assist the editing of pre-trained LLMs, there also exist branches of works that do not rely on external parameters or memory. Concretely, global optimization strategies aim to inject new knowledge into LLMs by updating all parameters, i.e., $\phi$ in Eq. (15). Through fine-tuning model parameters with specific designs to ensure the preservation of knowledge irrelevant to the target knowledge $t^{*}$, the LLMs are endowed with the ability to absorb new information without altering unedited knowledge. Generally, the goal of global optimization methods can be formulated as follows:
@@
where $f_{\phi}$ denotes the LLM before editing with the pre-trained parameter $\phi$, and $f_{\phi^{*}}$ denotes the edited LLM with updated parameter $\phi^{*}$. Generally, these methods focus more on the precision and generality of desirable knowledge, as the fine-tuning process ensures that the LLMs achieve satisfactory results regarding the edits and relevant knowledge. Nevertheless, as fine-tuning affects all parameters, they cannot easily preserve the locality of edited models, i.e., maintaining consistent output for unedited knowledge (Yao et al., 2023). In practice, directly applying fine-tuning strategies typically exhibits suboptimal performance on KME due to overfitting concerns (Wang et al., 2023c; Meng et al., 2023). Furthermore, fine-tuning large language models is also time-consuming and lacks scalability for multiple edits. Therefore, recently, motivated by these two challenges in fine-tuning, several global optimization works have been proposed and can be categorized as constrained fine-tuning methods and intermediate fine-tuning methods. Note that this section primarily focuses on methods from the model training perspective. Additionally, certain studies (Jiang et al., 2024; Gangadhar and Stratos, 2024) address the overfitting challenge by constructing more a comprehensive $\mathcal{X_{E}^{\prime}}$ with the following fine-tuning goal:
(20) $\min\mathbb{E}_{e\in\mathcal{E}}\mathbb{E}_{x,y^{*}\in\mathcal{X}_{e}^{\prime},{\mathcal{Y}^{*}_{e}}^{\prime}}\mathcal{L}(f_{\phi^{*}}(x),y^{*}),\ \text{where}\ \ f_{\phi^{*}}=M(f_{\phi};\mathcal{E}),$
[…]
@@ 5.3.2. Constrained Fine-tuning
The updates are regularized by restricting the norm of parameters before and after updating. RECT (Gu et al., 2024) adopts a similar yet simpler approach, specifically modifying only the top-kof parameters with the largest numerical updates during fine-tuning. Although restricting the norm is helpful in preventing the forgetting of original knowledge, the fine-tuning process can be less effective. To deal with this, RecAdam (Chen et al., 2020), in addition to the norm constraint, applies an annealing technique to control the ratio between the parameter norm and the fine-tuning loss as follows:
@@
Here $k$ and $t_{0}$ are hyper-parameters. $t$ is the number of fine-tuning steps. Such a design enables a gradual fine-tuning process that prevents massive parameter updates at the beginning. Motivated by the intuition of regularization to preserve original knowledge, PPA (Lee et al., 2022) employs LoRA (Hu et al., 2021) in the feed-forward (FFN) layers of the transformer decoder. LoRA is proposed to train the expansion/reduction matrix, instead of the model parameter $\phi$, to improve training speed by only updating parameters with a low intrinsic rank via dimensionality reduction. PPA leverages plug-in modules trained with constraints via LoRA to keep original knowledge intact. Moreover, the authors assess whether the content of the inputs falls within the scope of $\mathcal{X}_{\mathcal{E}}$ using the K-adapter module (Wang et al., 2020), and redirect such inputs to the new plug-in modules. This information is then used to determine whether to employ LoRA within the FFN layers. Furthermore, MELO (Yu et al., 2024) clusters the edits and employs multiple non-overlapping LoRA blocks for fine-tuning each cluster separately, thereby mitigating the issue of catastrophic forgetting. F-Learning (Forgetting before Learning) (Ni et al., 2023) proposes another approach to preserve original knowledge, which learns knowledge parameters $\Delta\phi$ that indicates old knowledge to be forgotten, defined as follows:
@@ 5.3.3. Intermediate Fine-tuning Strategies
While constrained fine-tuning techniques have demonstrated remarkable efficacy in a variety of NLP tasks (Wortsman et al., 2022; Ziegler et al., 2019; Bakker et al., 2022), they still exhibit instability and high computational cost when applied to KME, primarily due to the necessity of altering all parameters (Yao et al., 2023). A potential solution to address this challenge is to utilize an intermediate model to obtain the updated parameters in an efficient manner. Such an intermediate model is required to maintain significantly fewer parameters to ensure efficiency (Cheng et al., 2024). In general, recent works have widely adopted the Hyper-Network (Ha et al., 2016) as the intermediate model. Specifically, the Hyper-Network is a small network that generates the weights for a larger network, referred to as the main network. Specifically, the Hyper-Network takes inputs that contain information about the structure of the weights and generates the weights for layers in the main network. With the generated weights, the main network is updated to map input data to desired output targets. The updating process for the main network, denoted as $\phi$, can be defined as follows:
(26) $\phi^{*}$ $=\phi+\Delta\phi,\ \ \text{where}\ \ \Delta\phi=\text{H}(\nabla_{\phi}\mathcal{L}(f_{\phi}(x),y^{*}))\ \ \text{and}\ \ x,y^{*}\in\mathcal{X}_{\mathcal{E}},\mathcal{Y}^{*}_{\mathcal{E}},$
@@
However, hyper-networks are generally not capable of updating large language models due to the massive parameter size. To tackle this challenge, MEND (Mitchell et al., 2022a) adopts a mechanism referred to as gradient decomposition. In particular, it leverages small auxiliary editing networks to transform the gradients obtained by standard fine-tuning into edits of weights in a pre-trained model. As gradients are generally high-dimensional objects, a low-rank decomposition of the gradients is utilized to achieve the transformation. Particularly, MEND parameterizes the gradient mapping functions as MLPs with a single hidden layer, such that a significantly small number of parameters are required, compared with the edited models. In this manner, MEND enables fast model editing that can operate on considerably large pre-trained language models. Moreover, KGEditor (Cheng et al., 2024) proposes to combine the benefits of memory-based methods and hyper-networks to ensure flexibility and further reduce computation costs. Particularly, KGEditor introduces an additional layer with the same architecture of FFN layers for storing knowledge. Then it constructs a hyper-network based on a bi-directional LSTM (Hochreiter and Schmidhuber, 1997) that encodes embeddings of triples. In this manner, KGEditor becomes an efficient way to edit knowledge graph embeddings.
@@ 5.3.4. Summary
Global optimization methods typically apply specific fine-tuning restrictions to regularize parameter updates, namely constrained fine-tuning strategies. This is to prevent overfitting and ensure the model’s performance on the unedited knowledge. One crucial advantage of such strategies is its generality regarding the relevant knowledge, i.e., in-scope inputs $\mathcal{X}_{e}$ of edit $e$. As the global optimization affects all parameters in a language model, the relevant knowledge in it will also be edited, thereby generalizing to such knowledge. On the other hand, the high computation costs of fine-tuning all parameters also motivate researchers to propose intermediate fine-tuning strategies that leverage hyper-networks. Furthermore, global optimization methods are mostly model-agnostic, which means they can be applied to other editing methods. Nevertheless, such possibilities are less explored in the context of KME. In terms of the drawbacks, global optimization methods are suboptimal in maintaining the locality of edited models, as the optimization can easily influence unedited knowledge. Hence, it is crucial to achieve a balance between generality and locality when optimizing language models with specific constraints or intermediate designs.
@@ 5.4.1. Overview
In the following subsections, we first introduce the concept of knowledge neuron in LLMs, which are specific neurons that store factual knowledge and can be activated to generate the desirable answer based on a certain query $x$. Then we discuss two local modification strategies for KME: (1) the groundtruth-based strategies, which identify and edit knowledge neurons based on the supervision signal provided by the groundtruth; (2) the prompt-based strategies, which locate knowledge neurons based on the input prompts.
@@
(28) $\text{SelfAtt}_{i}(\textbf{x})=\text{Softmax}\left(\textbf{q}_{i}\textbf{k}_{i}^{\top}\right)\textbf{v}_{i},\quad\text{FFN}(\textbf{h})=\text{GELU}\left(\textbf{h}\textbf{W}_{1}\right)\textbf{W}_{2}.$
@@
5.4.2. Groundtruth-based Strategies
Based on the knowledge neuron view of the FFN layer weights in pre-trained LLMs, various groundtruth-based methods are proposed to locate and edit the pre-trained LLMs. Generally, they perform editing in a top-down manner, utilizing the supervision signal provided by the correct groundtruth $y^{*}$. As an exemplar work, KD (Dai et al., 2022) proposes to change each weight $w^{(l)}_{i}$ (i.e., the $i$-th weight in the $l$-th layer of FFN) from 0 to the pre-trained value $\hat{w}^{(l)}_{i}$ and calculates the cumulative change in the probability of predicting the output $y^{*}$ with input $x$, where the weights with a high cumulative probability are considered relevant for knowledge regarding $y^{*}$. DEPN (Wu et al., 2023) proposes a similar cumulative probability-based strategy to detect knowledge neurons that store privacy knowledge. In contrast to locating and editing an individual weight ${w}^{(l)}_{i}$, ROME (Meng et al., 2022) proposes to update an entire FFN layer to encode the new knowledge of $y^{*}$. Specifically, they view the second layer weights $\textbf{W}_{2}$ in the FFN layer in Eq. (28) as a linear associative memory (Kohonen, 1972; Anderson, 1972) in the form of $\textbf{K}\textbf{W}_{2}=\textbf{V}$, where the keys K and values V associated with $\textbf{W}_{2}$ can be directly calculated via pseudoinverse. With such a view of $\textbf{W}_{2}$ in the FFN layer, the optimization objective of updating it into $\hat{\textbf{W}}_{2}$ to encode new knowledge in the edit $e=(s,r,o\rightarrow o^{*})$ can be formulated as follows:
@@ 5.4.2. Groundtruth-based Strategies
Here $\textbf{k}^{*}$, which should encode the information of the subject $s$, is calculated by sampling multiple $x\sim\mathcal{X}_{e}$ and taking the average of the outputs from the first dense layer of the FFN. The target activation $\textbf{h}^{*}$ is calculated via optimizing the probability of outputting the correct answers $y^{*}\in\mathcal{Y}_{e}$ of the pre-trained LLM via the subsequent layers. Then, an efficient rank-one update is conducted on the weights $\textbf{W}_{2}$ according to Eq. (29), such that after the update, the edited FFN layer can output the correct hidden representation $\textbf{h}^{*}$ conducive to the generation of the right answer $y^{*}$ from $\textbf{k}^{*}$. The ROME framework has been shown to generalize to the large Mamba model (Sharma et al., 2024). Recently, MEMIT (Meng et al., 2023) proposes to further generalize the above editing strategy of the FFN layers of pre-trained LLMs to the mass editing of different knowledge. Particularly, with $u$ new edits $\{e_{1},e_{2},\dotsc,e_{u}\}$ that are required to be updated in the weights $\textbf{W}_{2}$, the mass knowledge editing problem can be formulated as the following optimization problem:
@@
where $\textbf{k}_{i}$, $\textbf{v}_{i}$ are the original key, value pairs associated with the weights $\textbf{W}_{2}$ (i.e., row vectors in matrices K, V in Eq. (29)), whereas $\textbf{k}_{i}^{*}$, $\textbf{v}^{*}_{i}$ are the updated key, value pairs calculated from the $i$-th edit $e_{i}$ as with Eq. (29). In addition, since multiple edits are required, the update is shared among different MLP layers, which is conducted in a top-down manner to prevent the potential issue of editing layers that could affect the ones that have already been edited. The residual for each edit is spread evenly over the range of the critical FFN layer. The strategy of residual attribution has recently been improved by PMET (Li et al., 2024c), which adopts a square root strategy to spread residuals to bottom FFN layers such that more precise information can be conveyed to critical layers. Furthermore, EMMET (Gupta et al., 2024) generalized ROME and MEMIT by formulating the mass knowledge editing problem as a preservation (of irrelevant knowledge)-memorization (of new knowledge) constrained optimization problem, where they derive closed form weight update formulae when the edit is exact, i.e., $\textbf{k}^{*}_{i}\hat{\textbf{W}}_{2}=\textbf{v}^{*}_{i}$ instead of minimizing the MSE in Eq. (30).
From the application’s perspective, to remove toxic knowledge of LLM, DINM (Wang et al., 2024d) identifies layers that store toxic knowledge with the discrepancy of toxic/non-toxic sequence embeddings, and uses the non-toxic samples to locally modify the weights of identified layers.